OpenStudy (ammarah):

Evaluat the function without a calculator sec 330 @kirbykirby

OpenStudy (kirbykirby):

Hm I would like to explain this more graphically to make it clear but my open study is starting to lag a lot too

OpenStudy (kirbykirby):

Actually let me try and see..I hope this works:

OpenStudy (kirbykirby):

 since you can write a coordinate on the unit circle as a pair (cos x, sin x) I hope you see that in the picture above, cos(330) = cos(30) because it is just the same x-coordinate and hopefully you remember that cos(30) = rad(3) / 2 so 1/cos(30) = 1/cos(330) = sec 330 = 1/(rad(3)/2) = 2/rad(3) = (2 rad(3))/3